OpenStudy (anonymous):

A hemispherical tank has a total surface area of 243pi ft^2 (this includes the circular base area) Find the volume Please help!!

OpenStudy (anonymous):

\[ (1)\ A_{hemisphere} = 2\pi r^2\] \[(1)\ V_{hemisphere} = \frac{2}{3}\pi r^3\] We use 1 to find r \[ 243\pi ft^2 = 2\pi r^2\] \[ 121.5 ft^2 = r^2 \ (divide\ by\ 2 \pi)\] \[ 11.0 ft = r \ (square root)\] So now that we have r, we plug it back into our volume equation: \[ V_{hemisphere} = \frac{2}{3}\pi r^3\] \[ V_{hemisphere} = \frac{2}{3}\pi (11.0ft)^3\] \[ V_{hemisphere} = \frac{2}{3}\pi \cdot 1331ft^3 \] \[ V_{hemisphere} = 887\pi ft^3 \]

OpenStudy (anonymous):

Thank you sooo much!!

OpenStudy (anonymous):

u knew it was a half circle with a line going to the middle right?

OpenStudy (anonymous):

A hemisphere is half a sphere, which is a 3D object. A semicircle is a 2D object. I assumed it was this: http://www.k6-geometric-shapes.com/image-files/hemisphere.jpg

OpenStudy (anonymous):

yea exactly! thank you!!